PPARC Proposal, 30 January, 2002

Athanasios Dedes

University of Bonn, Germany

Footprints from Supersymmetry at B-factories and Colliders

1. Rare B-decays 2. Supersymmetry 3. The footprints 4. The strategy

Footprints from Supersymmetry at B-factories and Colliders  PPARC proposal 1

Rare B-decays

• b-quark : an important laboratory for the exploration of the Standard Model and Beyond

• Large mass and Long Lifetime

• Large production cross section in e^{+}e^{−} and hadron collisions

• The strong force binds b-quarks and d or s-antiquarks making B_{d} or B_{s} mesons

• Huge data samples of B-meson decays → B-factories → CLEO, Belle, Babar

• Huge data samples of B-meson decays are also expected at Tevatron and LHC

• Rare B-decays

• They appear only at the Quantum Level in the SM

• Flavour Changing Neutral Currents (FCNC) b → s, d

Footprints from Supersymmetry at B-factories and Colliders  PPARC proposal 2

Rare B-decays

• Limits on some Rare B-decays

Type  Channel  Expt. Expt. value SM prediction
Lepton Flavour B_{d}^{0} → e^{±}μ^{∓}  CDF < 3.5 × 10^{−6}  0
B_{s}^{0} → e^{±}μ^{∓}  CDF < 6.1 × 10^{−6}  0
FCNC  B_{d}^{0} → μ^{+}μ^{−}  CDF < 8.6 × 10^{−7} 1.5 × 10^{−10}
B_{s}^{0} → μ^{+}μ^{−}  CDF < 2.0 × 10^{−6} 3.5 × 10^{−9}
B^{+} → K^{+}μ^{+}μ^{−} BaBar < 1.3 × 10^{−6} 0.4 × 10^{−6}
B^{0} → K^{∗0}μ^{+}μ^{−} Belle < 3.0 × 10^{−6}  1 × 10^{−6}
B^{0} → Kl^{+}l^{−} Belle 0.75 × 10^{−6} 0.35 × 10^{−6}
B^{0} → X_{s}γ  CLEO 3.2 × 10^{−4}  3.4 × 10^{−4}

√ The difference between the experimental bounds and the SM prediction is a “Window” for Physics Beyond the SM √ Example : B_{s} → μ^{+}μ^{−} could be soon found at Tevatron

Question

Can Supersymmetry be revealed in such Rare B-meson decays?

Footprints from Supersymmetry at B-factories and Colliders  PPARC proposal 3

Supersymmetry

• Supersymmetry unites particles of different spin

• has profound implications for fundamental physics

• can be enlarged to include gravity

• offers a hope of solving the hierarchy problem

• A minimal supersymmetric extension of the SM has been constructed, which is known as MSSM

• unifies strong and electroweak forces

• Gravity plays an essential role in supersymmetry breaking • offers a candidate particle to account for dark matter • predicts a light Higgs boson

• predicts new particles-partners of the SM ones

⟶ Observables are affected

• Example : In MSSM the B-decay B_{s} → μ^{+}μ^{−} and the muon anomalous magnetic moment are enhanced with respect to the SM

Footprints from Supersymmetry at B-factories and Colliders  PPARC proposal 4

The footprints

A. Dedes, H. K. Dreiner and U. Nierste, Phys. Rev. Lett. 87, 251804 (2001)

√ Striking correlation of the Rare decay B_{s} → μ^{+}μ^{−} and the muon anomalous magnetic moment √ More than 10 events can be seen at Tevatron/CDF √ Crucial test for the SO(10) Grand Unified Models

Footprints from Supersymmetry at B-factories and Colliders  PPARC proposal 5

The Strategy

• B-factories (CLEO, Babar, Belle)

– Define the Rare B-decays which are affected by SUSY – Calculate quantum corrections for these decays

• Colliders (Tevatron)

– Search for Rare B-decays at Tevatron

– Search for direct evidence of supersymmetry

– Make Monte-Carlo analysis

• B-factories and Colliders – Search for correlations between observables

– Define or restrict the unknown parameters of SUSY • Build a bridge between Cavendish and DAMTP!

– Apply to Grand Unified and/or to Superstring Models